. Assume the US can make 300 PCs if it uses all its resources or it can make 300 bushels of wheat. If the tradeoff is constant, calculate the opportunity cost of 1 PC. Then calculate the opportunity cost of 1 bushel of wheat.
Answer #1

Opportunity cost is the value of next best alternative foregone.

300 PCs = 300 bushels of wheat

1 PC = 300/300 bushel of wheat = 1 bushel of wheat

The opportunity cost of 1 PC is 1 bushel of wheat.

300 bushels of wheat = 300 PCs

1 bushel of wheat = 300/300 = 1 PC

The opportunity cost of 1 bushel of wheat is 1 PC.